# B3C: A Minimalist Approach to Offline Multi-Agent Reinforcement Learning

### v2

@@ 3.1 Revisiting Behavior Cloning: Over-regularization
3.2 Critic Clipping: Alleviate Overestimation
In order to simultaneously avoid unstable critic training and over-regularization, we propose a simple technique called Critic Clipping (CC) for training the centralized critic, which integrates with BC regularization for training policies. CC clips the critic target value when calculating the target value during policy evaluation, which can lead to overestimation based on the maximum return in the dataset. Unlike conservative value estimation, which requires additional sampling or modeling of the behavior policy and introduces underestimation bias, CC is easy to implement and proves effective by limiting overestimation to below a certain threshold. Here, the reason for using the maximum return rather than the average return as the clipping standard is as follows: The learned RL policy is expected to outperform the behavior policy, and therefore, the expected return, which the critic approximates, is expected to be larger than the average return in the dataset. For simplicity, we use the maximum return, representing the best episode in the dataset. To allow further flexibility, we can use a scaled value of the maximum return for the clipping value. Note that we observe using the maximum return is sufficient empirically, as we will discuss in Sec.4.3.2. The corresponding loss function of the critic is given by $\mathcal{L}_{CC}(\bm{\phi},\psi)=$
$\mathbb{E}_{(s,\bm{\tau},\bm{a},r,s^{\prime},\bm{\tau^{\prime}})\sim\mathcal{D}}\left[\Big{(}y^{jt}-Q_{jt}(s,\bm{\tau},\bm{a};\bm{\phi})\Big{)}^{2}\right],~{}~{}\mbox{where}$
$y^{jt}=r+\gamma~{}\underbrace{\mbox{Max}\Big{[}Q_{jt}(s^{\prime},\bm{\tau^{\prime}},\bm{\pi}(\bm{\tau^{\prime}};\bm{\theta^{-}});\bm{\phi^{-}}),~{}~{}R^{*}\Big{]}}_{\text{Critic Clipping}}$
$~{}~{}~{}~{}~{}~{}~{}\mbox{and}~{}~{}R^{*}=M\times\max_{d}\sum_{t=1}^{T}r_{d,t}.$ (4)
Here, $M$ is a scalar value, and $\max_{d}\sum_{t=1}^{T}r_{d,t}$ represents the maximum return in the dataset where $r_{d,t}$ is the reward at time-step $t$ of $d$-th episode in the dataset. For simplicity, we use $M=1$ in most cases during the experiments. We observed that this straightforward use of $M$ is sufficient in most cases, reducing the burden of hyperparameter tuning. We provide an ablation study on this value in Sec. 4.3.2.
3.3 B3C with Value Factorization
@@
$~{}~{}~{}~{}\mathbb{E}_{\mathcal{D}}\Big{[}-\bm{w}\underbrace{Q_{jt}(s,\bm{\tau},\pi^{1}(\tau^{1}),\cdots,\pi^{N}(\tau^{N}))}_{FACMAC}~{}~{}+$ (5)
$\beta\sum_{i=1}^{N}\underbrace{(\pi^{i}(\tau^{i})-a^{i})^{2}}_{BC}\Big{]},\mbox{where}~{}\bm{w}=\frac{\alpha}{\frac{1}{N}\sum|Q_{jt}(s,\bm{\tau},\bm{a})|},$
@@
$\bm{\pi}=(\pi^{1},\cdots,\pi^{N})$ and $\bm{a}=(a^{1},\cdots,a^{N})$ represent the joint policy and the joint action, respectively. Here, $\bm{\theta}_{\bm{\pi}}$ denotes the parameters of the joint policy, and $\mathcal{D}$ represents the given offline dataset. We normalize the Q-value to ensure robustness to scale, following Fujimoto & Gu (2021). Note that there are two important hyperparameters, $\alpha$ and $\beta$, which balance the trade-off between RL and BC. We will refer to $\alpha$ and $\beta$ as the RL coefficient and the BC coefficient, respectively. A key difference from Fujimoto & Gu (2021) is the introduction of the BC coefficient, $\beta$, alongside the RL coefficient, $\alpha$. For hyperparameter tuning, we recommend practitioners start by fixing $\beta=1$ and tuning $\alpha$, following Fujimoto & Gu (2021). If over-regularization persists, $\beta$ can then be adjusted, which has been shown to be effective in multi-agent particle environments.
@@
Empirical Observation regarding value factorization: We consider three value factorization methods discussed in Sec.2.1: non-mono, mono, and vdn. We empirically observe that non-mono outperforms mono in most cases, which differs from the findings in the online setting discussed in Sec. 2.1. We argue that limiting representational capacity, which is inherent to monotonic factorization, often leads to reduced performance in offline settings. We provide an ablation study for this in Sec. 4.3.3.
@@
As summarized above, we update the joint policy to maximize the critic with BC regularization by minimizing Eq. 5, and we update the centralized critic to minimize the TD error with clipping, as shown in Eq. 3.2. We refer to the proposed algorithm—FACMAC integrated with Behavior Cloning regularization and Critic Clipping—as FACMAC+B3C. Additionally, we also apply B3C to MA-TD3, referred to as MA-TD3+B3C.
4 Experiments
Table 1: Results of MA-TD3+B3C and the considered baselines, averaged over five seeds. Boldface numbers indicate the highest score or a comparable score among the algorithms.
Task-Dataset Avg R Max R OMAR CFCQL MADIFF MA-TD3+BC MA-TD3+BC (ours) MA-TD3+B3C (ours)
CN-e 100.0 163.1 114.9$\pm$2.6 112.0$\pm$4.0 95.0$\pm$5.3 108.3$\pm$3.3 100.6$\pm$8.3 102.3$\pm$6.7
CN-mr 9.5 123.2 37.9$\pm$12.3 52.2$\pm$9.6 30.3$\pm$2.5 15.4$\pm$5.6 53.8$\pm$4.8 53.8$\pm$4.8
CN-m 27.8 145.0 47.9$\pm$18.9 65.0$\pm$10.2 64.9$\pm$7.7 29.3$\pm$4.8 55.9$\pm$25.8 63.5$\pm$9.7
CN-r 0.0 103.7 6.9$\pm$3.1 62.2$\pm$8.1 6.9$\pm$3.1 9.8$\pm$4.9 72.6$\pm$7.0 73.3$\pm$6.6
PP-e 100.0 275.7 116.2$\pm$19.8 118.2$\pm$13.1 120.9$\pm$14.6 115.2$\pm$12.5 108.1$\pm$22.9 124.6$\pm$21.2
PP-mr 9.6 128.0 47.1$\pm$15.3 71.1$\pm$6.0 62.3$\pm$9.2 28.7$\pm$20.9 75.2$\pm$6.8 76.5$\pm$8.6
PP-m 48.9 196.2 66.7$\pm$23.2 68.5$\pm$21.8 77.2$\pm$10.4 65.1$\pm$29.5 91.3$\pm$35.7 104.3$\pm$11.9
PP-r 0.0 61.4 11.1$\pm$2.8 78.5$\pm$15.6 3.2$\pm$4.0 5.7$\pm$3.5 105.0$\pm$16.7 105.5$\pm$10.0
World-e 100.0 270.7 110.4$\pm$25.7 119.7$\pm$26.4 122.6$\pm$14.4 110.3$\pm$21.3 120.7$\pm$17.9 124.3$\pm$5.1
World-mr 12.0 138.7 42.9$\pm$19.5 73.4$\pm$23.2 57.1$\pm$10.7 17.4$\pm$8.1 75.2$\pm$23.1 75.8$\pm$10.9
World-m 58.1 206.8 74.6$\pm$11.5 93.8$\pm$31.8 123.5$\pm$4.5 73.4$\pm$9.3 119.3$\pm$18.2 119.7$\pm$16.7
World-r 0.0 63.1 5.9$\pm$5.2 68.0$\pm$20.8 2.0$\pm$3.0 2.8$\pm$5.5 98.3$\pm$49.7 101.0$\pm$16.8
4.1 Experimental Setup
@@
$(1)$ Three multi-agent particle environments (provided by Pan et al. (2022) and generated using MA-TD3 (Ackermann et al., 2019)): (a) Cooperative Navigation (CN), where three agents cover three landmarks without colliding, requiring coordination and teamwork; (b) Predator-Prey (PP), where three slower predators cooperate to capture a faster, pre-trained prey; and (c) World, where four predators attempt to catch two prey that aim to eat food while strategically using forest hiding spots for protection. For each task, evaluation is performed using four datasets of varying quality: expert, medium, medium-replay, and random.
$(2)$ Three fully observable multi-agent MuJoCo environments (provided in Wang et al. (2024) and generated using HAPPO (Kuba et al., 2021)): We consider three tasks including 3-Agent Hopper (HC), 2-Agent Ant (Ant), and 6-Agent HalfCheetah (HC). For each task, evaluation is performed using four datasets of varying quality: expert, medium, medium-replay, and medium-expert.
@@
Baselines.    We compare the proposed method against various baselines for each environment. For the multi-agent particle environments, we use MA-TD3+BC (Pan et al., 2022), OMAR (Pan et al., 2022), MADIFF (Zhu et al., 2023), and CFCQL (Shao et al., 2024) as baselines. For both multi-agent Mujoco environments, we use the multi-agent versions of BCQ (Fujimoto et al., 2019) and CQL (Kumar et al., 2020), as well as IQL (Yang et al., 2021), OMAR (Pan et al., 2022), CFCQL (Shao et al., 2024), and OMIGA (Wang et al., 2024).
Implementation and Hyperparameters.   To ensure a fair comparison, we implement the proposed method using the released CFCQL (Shao et al., 2024) code for the multi-agent particle environments and the released OMIGA (Wang et al., 2024) code for the multi-agent Mujoco environments. The RL and BC coefficients in Eq. 5 are critical hyperparameters in our approach. As mentioned in Sec. 3.3, we initially set $\beta=1$ and tune $\alpha$ for each task, as done in Fujimoto & Gu (2021). We observe that $\beta=1$ is sufficient for multi-agent Mujoco, but tuning $\beta$ improves performance in the multi-agent particle environments. For another hyperparameter, the scalar clipping value $M$ in Eq. 3.2, we use $M=1$ for simplicity, except for the Ant task with the medium-replay dataset, where a lower $M$ significantly improves performance. An ablation study on this value is provided in Sec. 4.3.2. The detailed hyperparameters used are listed in Appendix A.
Figure 2: Performance difference between the worst-performing seed of MA-TD3+BC and that of MA-TD3+B3C, with negative values indicating that MA-TD3+BC performs worse than MA-TD3+B3C. A value of $-5\%$ represents that BC performs $5\%$ worse than B3C.
4.2 Numerical Results
Refer to caption Refer to caption
(a) Ant2x4-return (b) Ant2x4-target
Refer to caption Refer to caption
(c) HC6x1-return (d) HC6x1-target
Figure 3: Test return and target value during policy evaluation in the training of BC (blue) and B3C (red). The black dotted line indicates the moment when the target value starts to diverge.
@@ 4.2 Numerical Results
Multi-agent particle environments.    The results are shown in Table 1. As discussed in Sec. 3.1, MA-TD3+BC from prior works suffers from over-regularization, which is evident from its performance on medium-replay and random datasets, where it significantly underperforms compared to the baselines. Thus, we first further tune the coefficients of RL and BC, which is referred to as MA-TD3+BC (ours) in Table 1. This simple tuned version substantially improves performance and even outperforms the baselines in some cases. However, this simple hyperparameter tuning is insufficient: BC regularization alone is often unstable. We claim that the proposed CC can address this instability. For this, Fig. 2 illustrates the performance difference between the worst-performing seed among 5 seeds of MA-TD3+BC and that of MA-TD3+B3C, where $-5\%$ indicates that MA-TD3+BC performs $5\%$ worse than MA-TD3+B3C. As shown in Fig. 2, using BC alone results in a significant performance drop in the worst-performing seed compared to B3C due to its instability, demonstrating that B3C, which incorporates critic clipping, stabilizes learning.
@@
Figure 4: Performance with respect to the RL coefficient $\alpha$ in the Halfcheetah environment.
Table 2: Results of FACMAC+B3C and the considered baselines across six tasks and several datasets in multi-agent Mujoco. Maximum and average returns are provided for each dataset. Dataset qualities are abbreviated as follows: ’expert’ as ’e,’ ’medium1’ as ’m1,’ and ’medium2’ as ’m2.’ Combinations of datasets are denoted as ’e-m1’ for ’e’ and ’m1,’ and ’m1-m2’ for ’m1’ and ’m2.’ Boldface numbers indicate the highest score or a comparable one among the algorithms.
Task-Dataset Avg R Max R BCQ-MA CQL-MA IQL OMAR CFCQL OMIGA FACMAC+B3C (ours)
Fully observable multi-agent MuJoCo provided in OMIGA (Wang et al., 2024) and generated by HAPPO (Kuba et al., 2021)
Hop-e 2452.0 3762.7 77.9 159.1 755.7 2.4 718.5 859.6 3619.7$\pm$1.6
Hop-m 723.6 2776.5 45.6 401.2 501.8 21.3 674.2 1189.3 3242.7$\pm$129.1
Hop-mr 746.4 2801.2 26.5 31.4 195.4 3.3 1380.2 774.2 736.8$\pm$469.4
Hop-me 1190.6 3762.7 54.3 64.8 355.4 1.4 383.0 709.0 3328.0$\pm$369.8
Ant-e 2055.1 2124.2 1317.7 1042.4 2050.0 313.5 1756.1 2055.5 2162.8$\pm$46.0
Ant-m 1418.7 1473.9 1059.6 533.9 1412.4 -1710.0 1159.6 1418.4 1516.5$\pm$14.8
Ant-mr 1029.5 1517.1 950.8 234.6 1016.7 -2014.2 1052.9 1105.1 1259.8$\pm$302.4
Ant-me 1736.9 2124.2 1020.9 800.2 1590.2 -2992.8 613.2 1720.3 2077.6$\pm$194.2
HC-e 2785.1 3866.1 2992.7 1189.5 2955.9 -206.7 4999.2 3383.6 5403.5$\pm$169.7
HC-m 1415.7 2113.5 2590.5 1011.4 2549.3 -265.7 4345.0 3608.1 4756.6$\pm$56.3
HC-mr 655.8 2132.6 -333.6 1998.7 1922.4 -235.4 3655.3 2504.7 4602.6$\pm$150.2
HC-me 2105.4 3866.1 3543.7 1194.2 2834.0 -253.8 5030.9 2948.5 5413.7$\pm$99.4
Partial observable multi-agent MuJoCo generated by ADER (Kim & Sung, 2023)
HC-k0-e 1394.3 1520.8 - - - 197.2 750.7 1390.1 1396.8$\pm$4.5
HC-k0-m1 1103.3 1332.9 - - - 189.6 443.7 1106.3 1141.6$\pm$18.9
HC-k0-m2 840.8 1157.4 - - - 839.1 766.5 847.6 1195.0$\pm$51.9
HC-k0-e-m1 1252.3 1520.8 - - - 136.4 542.6 1199.5 1307.1$\pm$27.3
HC-k0-e-m2 1121.7 1520.8 - - - 299.2 398.4 1097.9 1230.0$\pm$40.5
HC-k0-m1-m2 976.2 1332.9 - - - 709.1 1186.1 1027.3 1210.1$\pm$22.8
HC-k1-e 3766.0 3863.3 - - - 3232.6 3390.2 3760.8 3760.5$\pm$24.2
HC-k1-m1 1976.2 2350.0 - - - 2312.4 2356.0 2017.3 2508.1$\pm$55.7
HC-k1-m2 1223.9 1758.4 - - - 1108.9 1440.6 1196.5 2187.8$\pm$66.7
HC-k1-e-m1 2873.2 3863.3 - - - 2296.4 1949.6 2112.3 2682.0$\pm$175.4
HC-k1-e-m2 2486.0 3863.3 - - - 524.8 864.7 1088.2 1102.8$\pm$349.8
HC-k1-m1-m2 1591.6 2237.2 - - - 1410.6 1862.2 1633.5 $2$.6$\pm$84.1
Sw-e 430.9 438.9 - - - 395.3 403.3 430.7 430.3$\pm$2.6
Sw-m1 290.4 306.3 - - - 268.4 277.1 288.3 289.5$\pm$1.1
Sw-m2 142.7 158.5 - - - 97.9 130.0 142.6 155.3$\pm$1.7
Sw-e-m1 360.6 438.9 - - - 216.5 292.1 261.4 297.1$\pm$53.8
Sw-e-m2 286.8 438.9 - - - 113.8 138.6 148.2 211.9$\pm$8.4
Sw-m1-m2 216.3 306.3 - - - 226.4 222.7 205.8 189.9$\pm$7.7
Multi-agent Mujoco.   In this environment, we evaluate FACMAC+B3C, which incorporates non-monotonic value factorization with B3C. For fully observable cases, we use the performance results of BCQ-MA, CQL-MA, IQL, OMAR, and OMIGA as reported in  (Wang et al., 2024), while we generate results for CFCQL and FACMAC+B3C. For partially observable cases, we generate all results using their public code. As shown in Table 2, the proposed method, FACMAC+B3C, outperforms the baselines in most cases. Notably, across all datasets for the HalfCheetah environment and several datasets for the Hopper and Ant environments, FACMAC+B3C even outperforms the maximum return of the datasets.

### v3

@@ 3.1. Revisiting Behavior Cloning: Over-regularization
3.2. Critic Clipping: Alleviate Overestimation
In order to simultaneously avoid unstable critic training and over-regularization, we propose a simple yet effective technique called Critic Clipping (CC) for training the centralized critic, which integrates with BC regularization for training policies.
CC clips the critic target value when calculating the target value during policy evaluation, which can lead to overestimation based on the maximum return in the dataset. Unlike conservative value estimation, which requires additional sampling or modeling of the behavior policy and introduces underestimation bias, CC is easy to implement and proves effective by limiting overestimation to below a certain threshold. Intuitively, this prevents the critic from propagating excessively optimistic targets that would otherwise destabilize value learning across agents. Here, the reason for using the maximum return rather than the average return as the clipping standard is as follows: The learned RL policy is expected to outperform the behavior policy, and therefore, the expected return, which the critic approximates, is expected to be larger than the average return in the dataset. For simplicity, we use the maximum return, representing the best episode in the dataset. To allow further flexibility, we can use a scaled value of the maximum return for the clipping value. Note that we observe using the maximum return is sufficient empirically, as we will discuss in Sec.4.3.2. The corresponding loss function of the critic is given by
$\mathcal{L}_{CC}(\bm{\phi},\psi)=\mathbb{E}_{(s,\bm{\tau},\bm{a},r,s^{\prime},\bm{\tau^{\prime}})\sim\mathcal{D}}\left[\Big(y^{jt}-Q_{jt}(s,\bm{\tau},\bm{a};\bm{\phi})\Big)^{2}\right],\penalty 10000\ \penalty 10000\ \mbox{where}$
$y^{jt}=r+\gamma\penalty 10000\ \underbrace{\mbox{Min}\Big[Q_{jt}(s^{\prime},\bm{\tau^{\prime}},\bm{\pi}(\bm{\tau^{\prime}};\bm{\theta^{-}});\bm{\phi^{-}}),\penalty 10000\ \penalty 10000\ R^{*}\Big]}_{\text{Critic Clipping}}\penalty 10000\ \mbox{and}$ (5)
$R^{*}=M\times\max_{d}\sum_{t=1}^{T}r_{d,t}$. Here, $M$ is a scalar value, and $\max_{d}\sum_{t=1}^{T}r_{d,t}$ represents the maximum return in the dataset where $r_{d,t}$ is the reward at time-step $t$ of $d$-th episode in the dataset. For simplicity, we use $M=1$ in most cases during the experiments. We observed that this straightforward choice is sufficient in most cases, aligning with our minimalist design philosophy and reducing the burden of hyperparameter tuning. Although $M$ can be fine-tuned (e.g., smaller $M$ performed slightly better in one environment; see Appendix A), we did not observe a clear or consistent improvement from adjusting it. Additionally, this clipping does not constrain the learned policy to achieve returns below the maximum return in the dataset. The critic value provides an approximation of policy performance but does not constitute a hard upper bound, as its estimates are used for optimization guidance rather than to determine the final achievable return. We provide an ablation study on this in Sec. 4.3.2.
3.3. B3C with Value Factorization
@@
$\mathbb{E}_{\mathcal{D}}\Big[-\bm{w}$ $\underbrace{Q_{jt}(s,\bm{\tau},\pi^{1}(\tau^{1}),\cdots,\pi^{N}(\tau^{N}))}_{FACMAC}+\beta\sum_{i=1}^{N}\underbrace{(\pi^{i}(\tau^{i})-a^{i})^{2}}_{BC}\Big]$
$\mbox{where}\penalty 10000\ \bm{w}=\frac{\alpha}{\frac{1}{N}\sum|Q_{jt}(s,\bm{\tau},\bm{a})|},$ (6)
@@
$\bm{\pi}=(\pi^{1},\cdots,\pi^{N})$ and $\bm{a}=(a^{1},\cdots,a^{N})$ represent the joint policy and the joint action, respectively. Here, $\bm{\theta}_{\bm{\pi}}$ denotes the parameters of the joint policy, and $\mathcal{D}$ represents the given offline dataset. We normalize the Q-value based on the sample batch to ensure robustness to scale, following fujimoto2021minimalist. Note that there are two important hyperparameters, $\alpha$ and $\beta$, which balance the trade-off between RL and BC. We will refer to $\alpha$ and $\beta$ as the RL coefficient and the BC coefficient, respectively. A key difference from fujimoto2021minimalist is the introduction of the BC coefficient, $\beta$, alongside the RL coefficient, $\alpha$. This decouples the relative RL-BC weighting and the overall objective scale, enabling independent control of regularization strength, which can affect optimization in deep networks. For hyperparameter tuning, we recommend practitioners start by fixing $\beta=1$ and tuning $\alpha$, following fujimoto2021minimalist. If over-regularization persists, $\beta$ can then be adjusted, which has been shown to be effective in multi-agent particle environments.
@@
Empirical Observation regarding value factorization: We consider three value factorization methods discussed in Sec.2.1: non-mono, mono, and vdn. We empirically observe that non-mono outperforms mono in most cases, which differs from the findings in the online setting discussed in Sec. 2.1. We argue that limiting representational capacity, which is inherent to monotonic factorization, leads to reduced performance in offline settings. We provide an ablation study for this in Sec. 4.3.3.
@@
As summarized above, we update the joint policy to maximize the critic with BC regularization by minimizing Eq. 3.3, and we update the centralized critic to minimize the TD error with clipping, as shown in Eq. 3.2. We refer to the proposed algorithm—FACMAC integrated with Behavior Cloning regularization and Critic Clipping—as FACMAC+B3C. Additionally, we also apply B3C to MA-TD3, referred to as MA-TD3+B3C.
Table 1. Performance of MA-TD3+B3C and baselines on multi-agent particle environments, averaged over five seeds. Tasks include Cooperative Navigation (CN), Predator-Prey (PP), and World, each evaluated on expert (e), medium (m), medium-replay (mr), and random (r) datasets. “Avg R” and “Max R” denote the dataset average and maximum return. Boldface indicates the best or statistically comparable performance.
Task-Dataset Avg R Max R OMAR CFCQL MADIFF TD3+BC TD3+BC (ours) TD3+B3C (ours)
[…]
4. Experiments
4.1. Experimental Setup
@@
$(1)$ Three multi-agent particle environments (provided by pan2022plan and generated using MA-TD3 ackermann2019reducing): (a) Cooperative Navigation (CN), where three agents cover three landmarks without colliding, requiring coordination and teamwork; (b) Predator-Prey (PP), where three slower predators cooperate to capture a faster, pre-trained prey; and (c) World, where four predators attempt to catch two prey that aim to eat food while strategically using forest hiding spots for protection. For each task, evaluation is performed using four datasets of varying quality: expert, medium, medium-replay, and random.
$(2)$ Three fully observable multi-agent MuJoCo environments (provided in wang2024offline and generated using HAPPO kuba2021trust): We consider three tasks including 3-Agent Hopper (HC), 2-Agent Ant (Ant), and 6-Agent HalfCheetah (HC). For each task, evaluation is performed using four datasets of varying quality: expert, medium, medium-replay, and medium-expert.
$(3)$ Two partially observable multi-agent MuJoCo environments: 6-Agent HalfCheetah and 5-Agent Swimmer, each with differing degrees of partial observability and performance levels. We generated the offline datasets using ADER kim2023adaptive with varying levels of performance: expert, medium-1, medium-2, and their combinations. In both environments, each agent can observe its $K$ nearest neighbors. Here, $K$ determines the degree of partial observability. We consider $K=0,1$ for HC, which is denoted as HC-kKK, and $K=0$ for Swimmer (SW).
Baselines. We compare the proposed method against various baselines for each environment. For the multi-agent particle environments, we use MA-TD3+BC pan2022plan, OMAR pan2022plan, MADIFF zhu2023madiff, and CFCQL shao2024counterfactual as baselines. For multi-agent Mujoco environments, we use OMAR pan2022plan, CFCQL shao2024counterfactual, and OMIGA wang2024offline. In addition, we include TD3+BC, FACMAC+BC, and TD3+B3C to analyze the effect of value factorization and CC.
Implementation and Hyperparameters. To ensure a fair comparison, we implement the proposed method using the released CFCQL shao2024counterfactual code for the multi-agent particle environments and the released OMIGA wang2024offline code for the multi-agent Mujoco environments. The RL and BC coefficients in Eq. 3.3 are critical hyperparameters in our approach. As mentioned in Sec. 3.3, we initially set $\beta=1$ and tune $\alpha$ for each task, as done in fujimoto2021minimalist. We observe that $\beta=1$ is sufficient for multi-agent Mujoco, but tuning $\beta$ improves performance in the multi-agent particle environments. For another hyperparameter, the scalar clipping value $M$ in Eq. 3.2, we use $M=1$ for simplicity, except for the Ant task with the medium-replay dataset, where a lower $M$ significantly improves performance. An ablation study on this value is provided in Sec. 4.3.2. The detailed hyperparameters used are listed in Appendix A.
Figure 2. Worst-seed performance gap between MA-TD3+BC and MA-TD3+B3C. Negative values indicate that MA-TD3+BC performs worse (i.e., a gap of $5\%$ means its worst-seed performance is $5\%$ lower than that of MA-TD3+B3C).
4.2. Numerical Results
Table 2. Results of FACMAC+B3C and baselines on six Mujoco tasks across multiple datasets. Both maximum and average returns are reported. Dataset types are abbreviated as: expert (e), medium1 (m1), and medium2 (m2). Combined datasets are denoted as e-m1 and m1-m2. Bold indicates the best or comparable performance.
Task-Dataset Avg R Max R OMAR CFCQL OMIGA TD3+BC FACMAC+BC TD3+B3C FACMAC+B3C
Fully observable multi-agent MuJoCo provided in OMIGA wang2024offline and generated by HAPPO kuba2021trust
@@ 4.2. Numerical Results
[…]
Partial observable multi-agent MuJoCo generated by ADER kim2023adaptive
[…]
@@
Multi-agent particle environments. The results are shown in Table 1. As discussed in Sec. 3.1, MA-TD3+BC from prior works suffers from over-regularization, which is evident from its performance on medium-replay and random datasets, where it significantly underperforms compared to the baselines. Thus, we first further tune the coefficients of RL and BC, which is referred to as MA-TD3+BC (ours) in Table 1. This simple tuned version substantially improves performance and even outperforms the baselines in some cases. However, this simple hyperparameter tuning is insufficient: BC regularization alone is often unstable. We claim that the proposed CC can address this instability. Fig. 2 illustrates the performance difference between the worst-performing seed among 5 seeds of MA-TD3+BC and that of MA-TD3+B3C, where $5\%$ indicates that the worst-performing seed of MA-TD3+BC performs $5\%$ worse than MA-TD3+B3C. As shown in Fig. 2, using BC alone results in a significant performance drop in the worst-performing seed compared to B3C due to its instability, demonstrating that B3C, which incorporates critic clipping, stabilizes learning. Additionally, in terms of average performance, TD3+B3C outperforms the baselines in most cases. In particular, on the medium-replay and random datasets, which are generated by multiple behavior policies, TD3+B3C achieves state-of-the-art performance with large margins.
Multi-agent Mujoco. In this environment, we evaluate FACMAC+B3C, which combines non-monotonic value factorization with B3C. For fully observable cases, we use the reported results of OMAR and OMIGA wang2024offline, and generate results for TD3+BC, FACMAC+BC, and TD3+B3C for comparison. As shown in Table 2, FACMAC+B3C achieves the best overall performance across environments, often exceeding the dataset’s maximum return, particularly on HalfCheetah and several Hopper and Ant tasks. Moreover, unlike the baselines, FACMAC+B3C exhibits no performance collapse across any dataset, indicating stable and consistent learning.
The consistent improvement of FACMAC+B3C over FACMAC+BC, as well as TD3+B3C over TD3+BC, demonstrates the effectiveness of Critic Clipping in stabilizing value estimation and improving overall policy quality. This result highlights that the benefit of B3C is not limited to a specific architecture but generalizes across both factored and non-factored critics.
Furthermore, FACMAC+BC generally outperforms TD3+BC except for a few fully observable tasks such as Ant-me and HC-e, which aligns with findings from online MARL literature where value factorization improves coordination and representation efficiency. These results collectively confirm that integrating value factorization with B3C provides both high and stable performance in complex multi-agent continuous control settings.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Ant2x4-return (b) Ant2x4-target (c) HC6x1-return (d) HC6x1-target
Figure 3. Test return and target value during policy evaluation in the training of BC (blue) and B3C (red). The black dotted line indicates the moment when the target value starts to diverge.
@@
Figure 4. Performance with respect to the RL coefficient $\alpha$ in the partially observable Halfcheetah environment. Larger $\alpha$ increases the weight on the RL objective relative to BC regularization.
Method PP (N=3) PP (N=6) PP (N=9)
FACMAC+BC 71.1 $\pm$ 3.6 74.8 $\pm$ 20.1 85.6 $\pm$ 47.8
FACMAC+B3C 97.1 $\pm$ 6.0 90.8 $\pm$ 12.4 104.5 $\pm$ 16.8
Figure 5. Performance and target Q-values of FACMAC with BC and B3C for varying NN.
4.3. Analysis
4.3.1. Critic Clipping: B3C versus BC
Critic clipping alleviates overestimation by directly constraining the target value in policy evaluation. As mentioned earlier, BC regularization alone often makes learning unstable, and even the critic diverges. We have already shown instability problem in Sec. 4.2 through the worst performance of MA-TD3+BC in the multi-agent particle environments. Here, we provide further analysis of how critic clipping addresses this issue for better understanding in multi-agent Mujoco environments.
Fig. 3 presents the test return and the target value, $y^{jt}$ from Eq. 3.2, during the training of FACMAC+BC and FACMAC+B3C. The results indicate that FACMAC+BC, which relies solely on BC regularization, results in diverging value estimates. When this divergence occurs, marked by the dotted line in Fig. 3, the performance correspondingly degrades. However, B3C effectively mitigates this divergence—it occasionally tends to overestimate but either stabilizes quickly or avoids divergence entirely, ensuring more consistent performance. Note that this phenomenon does not always occur but is more likely when the RL objective is weighted more heavily than BC regularization; however, critic clipping can prevent it and enable stable learning.
Additionally, we evaluate FACMAC+BC and FACMAC+B3C in a partially observable Predator-Prey environment peng2021facmac with 3, 6, and 9 agents. This experiment aims to examine how the degree of overestimation changes as the number of agents $N$ increases. For a fair comparison, we normalize each dataset return by its average to compare target Q-values across the 3-, 6-, and 9-agent settings. As shown in Fig. 5, the target Q-values of FACMAC+BC diverge more rapidly as $N$ increases, indicating that overestimation becomes more severe with a larger number of agents due to the exponentially growing joint action space. In contrast, critic clipping in FACMAC+B3C effectively suppresses this divergence, keeping target values bounded even under increased coordination complexity. This stability corresponds to higher and more consistent performance, as shown in the table below, indicating that B3C maintains robustness as the number of agents increases and alleviates the instability induced by large-scale coordination in multi-agent offline RL.